A wire of 2ohm is redrawn so that its length becomes four times the resistance of the withdtawn wire is.? 

Grade:12

1 Answers

Susmita
5 years ago
Volume remains constant.
V'=V
Or,A'*4L=A*L
Or,A/A'=4
Resistance R=rho*L/A
R'/R=(L'/L)*(A/A')2=4*4=16
R'=16R=16*2=32 ohm